Consider the titration of a 33.0 mL sample of 0.180 M HBr with 0.210 M KOH. Determine each of the following:

Part A: the initial pH (Express your answer using three decimal places)

Part B: the volume of added base required to reach the equivalence point (Express your answer in millimeters)

Part C: the pH at 11.0 mL of added base (Express your answer using three decimal places)

Part D: the pH at the equivalence point (Express your answer as a whole number)
Answer #1

A)when 0.0 mL of KOH is added

Given:

M(HBr) = 0.18 M

V(HBr) = 33 mL

M(KOH) = 0.21 M

V(KOH) = 0 mL

mol(HBr) = M(HBr) * V(HBr)

mol(HBr) = 0.18 M * 33 mL = 5.94 mmol

mol(KOH) = M(KOH) * V(KOH)

mol(KOH) = 0.21 M * 0 mL = 0 mmol

We have:

mol(HBr) = 5.94 mmol

mol(KOH) = 0 mmol

0 mmol of both will react

remaining mol of HBr = 5.94 mmol

Total volume = 33.0 mL

[H+]= mol of acid remaining / volume

[H+] = 5.94 mmol/33.0 mL

= 0.18 M

use:

pH = -log [H+]

= -log (0.18)

= 0.7447

Answer: 0.745

B)

Balanced chemical equation is:

HBr + KOH ---> KBr + H2O

Here:

M(HBr)=0.18 M

M(KOH)=0.21 M

V(HBr)=33.0 mL

According to balanced reaction:

1*number of mol of HBr =1*number of mol of KOH

1*M(HBr)*V(HBr) =1*M(KOH)*V(KOH)

1*0.18 M *33.0 mL = 1*0.21M *V(KOH)

V(KOH) = 28.2857 mL

Answer: 28.3 mL

C)when 11.0 mL of KOH is added

Given:

M(HBr) = 0.18 M

V(HBr) = 33 mL

M(KOH) = 0.21 M

V(KOH) = 11 mL

mol(HBr) = M(HBr) * V(HBr)

mol(HBr) = 0.18 M * 33 mL = 5.94 mmol

mol(KOH) = M(KOH) * V(KOH)

mol(KOH) = 0.21 M * 11 mL = 2.31 mmol

We have:

mol(HBr) = 5.94 mmol

mol(KOH) = 2.31 mmol

2.31 mmol of both will react

remaining mol of HBr = 3.63 mmol

Total volume = 44.0 mL

[H+]= mol of acid remaining / volume

[H+] = 3.63 mmol/44.0 mL

= 8.25*10^-2 M

use:

pH = -log [H+]

= -log (8.25*10^-2)

= 1.0835

Answer: 1.083

D)

This is titration of strong acid and strong base.

At equivalence point, the solution would be neutral and hence pH would be 7.00

Answer: 7